Golden Dragon Restaurant obtained a $9400 loan at 8% compounded annually to replace some kitchen equipment. Prepare a complete amortization schedule if payments of $1840 (except for a smaller final payment) are made semiannually. (Do not round intermediate calculations. Round your answers to 2 decimal places. Leave no cells blank - be certain to enter "0" wherever required.
Answer #1
Answer:
Amortization schedule:
a b = a*4% c d = c-b a-d
Payments Opening balance Interest @4% Semi annual Payment Principal payment Closing balance
1 $9,400 376.00 $1,840.00 $1,464.00 $7,936.00
2 $7,936 317.44 $1,840.00 $1,522.56 $6,413.44
3 $6,413 256.54 $1,840.00 $1,583.46 $4,829.98
4 $4,830 193.20 $1,840.00 $1,646.80 $3,183.18
5 $3,183 127.33 $1,840.00 $1,712.67 $1,470.50
6 $1,471 58.82 $1,840.00 $1,781.18 ($310.68)
Total 1329.32 $11,040.00 $9,710.68